OpenStudy (anonymous):

Find the prime factorization of the number 117.

OpenStudy (anonymous):

117 = 3* 39 117 = 3 * 3 * 13 117 = 3^2 * 13

OpenStudy (anonymous):

is it 24?

OpenStudy (regan):

3 3 13